OpenStudy (awkwardpanda):

What is the solution to the equation? √2x+10 -6=2

OpenStudy (studygurl14):

Is it this: \[\sqrt{2x} + 10 - 6 = 2\] \[\sqrt{2x + 10} - 6 = 2\] Or \[\sqrt{2x + 10 -6}\] ?

OpenStudy (studygurl14):

sorry, add =2 at the end of the third option

OpenStudy (awkwardpanda):

the second one, its fine (:

OpenStudy (studygurl14):

Okay, so first, move the 6 to the other side. \[\sqrt{2x + 10}=2+6\] \[\sqrt{2x + 10}=10\] Now what?

OpenStudy (awkwardpanda):

next divide?

OpenStudy (studygurl14):

nope. Next, get rid of the square root. Do you know how to do that?

OpenStudy (awkwardpanda):

No.. ;-;

OpenStudy (studygurl14):

Well let's see. What's the opposite of taking the SQUARE root of a number? The power of two, or SQUAREing a number. So how would you cancel the square root? By squaring it. \[(\sqrt{2x + 10})^{2} =10^{2}\]

OpenStudy (studygurl14):

what next?

OpenStudy (awkwardpanda):

multiplying?

OpenStudy (awkwardpanda):

multiply 10^2 which would make it 100. right?

OpenStudy (radar):

Assuming that it is indeed the third option.\[\sqrt{2x+10-6}=2\]\[\sqrt{2x+4}=2\]2x+4=4 x=0 You did not really mean it was the third option, did you?

OpenStudy (radar):

It is so important to clearly represent the problem.

OpenStudy (awkwardpanda):

\[\sqrt{2x+10} -6=2\]

OpenStudy (radar):

Sorry, I read the StudyGurl14 post and thought it was yours. \[\sqrt{2x+10}=8\]\[2x+10 = 64\]After squaring both sides. 2x=54 after subtracting 10 from both sides x=27 after dividing both sides by 2

OpenStudy (radar):

Do you have that (x=27) as a possible solution?

OpenStudy (radar):

Good luck with these. One thing to be aware of, when a radical is squared there is a possibility of losing a sign.

OpenStudy (studygurl14):

oh stupid me. I did 2 + 6 as 10, not 8

OpenStudy (studygurl14):

sometimes I don't know where my mind is

OpenStudy (radar):

Anything but stupid, call it a typo, or a momentary lapse. You deserve a medal.

OpenStudy (studygurl14):

Thanks. Did you get the answer?

OpenStudy (studygurl14):

The "Did you get the answer?" was directed to @akwardpanda. The Thanks was directed to @radar.

OpenStudy (awkwardpanda):

Yes I did (: thank you both.